1mol of H2, 2.0mol of I2, 3mol of HI are injected into 1 Litre flask. What will be the equilibrium concentration of H2, I2, HI at 490^C? The equilibrium constant for the reaction is 45.9.
Answer #1

Ans Kc - 45.9 H₂ Iz 2 HI - 2 3 initial mole at eluilibrium |-X 3 + 2x 2-x Kc = [HI]2 [H22 Iz usog = (3+2x)? (1-x) (2-x) X = 0